OpenStudy (anonymous):

What is the value of the expression 7 + 52 × (7 - 3)

OpenStudy (cloverracer):

236

OpenStudy (anonymous):

Parenthesis, Exponants, Multiplication, Division, Addition, Subtraction PEMDAS First solve in the parenthesis 7-3 = 4

OpenStudy (anonymous):

Next comes multiplication. Multiply 52 by 4 = 208

OpenStudy (anonymous):

7-3 = 4 Then 52* 4 = 208 + 7 = 215

OpenStudy (anonymous):

Finally comes addition. 7 + 208 = 215 That is your answer

OpenStudy (anonymous):

@CloverRacer I'm not sure where they got 236. They must have done PEMDAS wrong.

OpenStudy (cloverracer):

7 + 52 * (7-3) 7 + 52 * 4 7 + 208 215 uh oh! careless mistake, my bad.